Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 01

1. Section 2.2: 41, 77, 87 (find the limit and need not plot it).

2. Section 2.3: problems 35, 49, 53.

3. Section 2.3: Read the ε − δ definition of ’Showing L is not a limit’ in p84 and review the decomposition steps shown in class that leads to this definition. Then do problem 57(b).